Job Costs At the end of August, Ellis Company had completed Jobs 40 and 42. Job 40 is for 200 units, and Job 42 is for 1,000 units. The following data relate to these two jobs: On August 4, raw materials were requisitioned for production as follows: 500 units for Job 40 at $18 per unit and 1,000 units for Job 42 at $16 per unit. During August, Ellis Company accumulated 400 hours of direct labor costs on Job 40 and 500 hours on Job 42. The total direct labor was incurred at a rate of $10 per direct labor hour for Job 40 and $18 per direct labor hour for Job 42. The predetermined factory overhead rate is $11.00 per direct labor hour. a. Determine the balance on the job cost sheets for Jobs 40 and 42 at the end of August. Job 40 $ Job 42 $ b. Determine the cost per unit for Jobs 40 and 42 at the end of August. If required, round your answers to the nearest cent. Job 40 $ Job 42 $
Answer #1

a) Job Cost sheet

Job 40 Job 42
Direct material 500*18 = 9000 1000*16 = 16000
Direct labor 400*10 = 4000 500*18 = 9000
Overhead 400*11 = 4400 5500
Total Cost 17400 30500

b) Cost per unit

Job 40 Job 42
Total Cost 17400 30500
Unit 200 1000
Cost per unit 87 30.50